Suppose a risk-free asset has a 5 percent return and a second asset has an expected return of 13 percent with a standard deviation of 23 percent. A portfolio consisting 10 percent of the risk-free asset and 90 percent of the second asset. What is the Sharpe ratio of this portfolio?
Answer #1

Expected return of Portfolio = 0.10(0.05) + 0.90(0.13) = 12.20%

Standard Deviation = 0.90(0.23) = 20.70%

Sharpe Ratio = (0.1220 - 0.05)/0.2070

Sharpe Ratio = 0.35